# Field-induced states and thermodynamics of the frustrated Heisenberg antiferromagnet on a square lattice

### v1

@@ III.2.1 Specific heat
The first case $h/J_{1}=4$ lies just below the saturation field that is $h_{\rm sat}/J_{1}=4.2$ for $J_{2}/J_{1}=0.55$ (see Eq. (7)). The result for the $N=36$ system is shown in Fig. 11(a) and serves as a benchmark. The proximity to saturation permits us to perform exact diagonalization (ED) that covers the complete spectrum up to sufficiently high temperatures to yield essentially exact results for the entire temperature range shown. This ED curve is free from the statistical errors that are inherent to FTLM and thus serves as a reference. We see that the FTLM curve is very close to the ED one, and also the TPQ curve deviates from these two only by an amount that is consistent with the estimated error. The features observed in Fig. 11(a) around $T/J_{1}=0.01$ and $0.05$ are most likely due to finite-size effects reflecting the finite system size $N=36$ and thus have no physical meaning.
@@
According to Subec. III.1.2, there should be no $m=1/2$ plateau at $T=0$ for $J_{2}/J_{1}=0.5$. If the case $s=1/2$ is analogous to the classical case shown in Fig. 2, the maximum of $C$ at $T/J_{1}\approx 0.1$ for $1\lesssim h/J_{1}\gtrsim 2$ observed in Fig. 12(a) might nevertheless indicate a transition out of an uuuduuud phase stabilized by thermal fluctuations. For $J_{2}/J_{1}=0.55$, we find an $m=1/2$ plateau in the range of fields $2.24\lesssim h/J_{1}\lesssim 2.60$, see Fig. 8(c). The maximum of $C$ for $T/J_{1}\lesssim 0.1$ that one observes in this range of fields in Fig. 12(b) probably corresponds to a transition out of this plateau state, as we have discussed in the context of Fig. 11(b). Invoking again the analogy with the classical case Fig. 2, the continuation of this maximum to lower fields $1.2\lesssim h/J_{1}\lesssim 2.2$ and at $T/J_{1}\lesssim 0.08$ might again correspond to a transition out of an uuuduuud state stabilized by thermal fluctuations. This suggests an intriguing scenario where the crossover from stabilization by quantum to thermal fluctuations around $h/J_{1}\approx 0.55$ would be accompanied by a minimum in the transition temperature.
In the low-field limit, the specific heat of Fig. 12(a,b) exhibits a maximum around $T/J_{1}\approx 0.2$. It is an interesting question whether this is just a fluctuation maximum or a signature related to non-trivial ground states, but this question goes beyond the scope of the present work.
@@ III.2.2 Entropy and magnetocaloric effect
Geometric frustration is expected to enhance the magnetocaloric effect [131, 132, 133, 130]. In the present case, we expect such an enhancement specifically at $J_{2}/J_{1}=1/2$ where the single-magnon spectrum becomes degenerate, recall Eq. (9). Such an enhancement was indeed observed in previous works [85, 86, 90] that scanned the ratio $J_{2}/J_{1}$ at a fixed temperature. Here, such an enhancement is observed as a pronounced minimum in the lowest three isentropes $T(h)$ around $h_{\rm sat}$ in Fig. 12(d).
@@ IV Conclusion
We have revisited the behavior of the $J_{1}$-$J_{2}$ FSAFM in a magnetic field, pushing previous investigations [72, 73, 74] to bigger system sizes and higher accuracy.
@@
Then we turned to the spin-1/2 case. First, we presented zero-temperature magnetization curves for several ratios of $J_{2}/J_{1}$ and provided evidence for the emergence of an $m=1/2$ plateau that can be understood as an uuuduuud state stabilized by quantum fluctuations. We find a relatively large stabilization regime $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$ that is shifted to values $J_{2}>J_{1}/2$ such that analytic treatment of this uuuduuud $m=1/2$ plateau requires expansion around a classically unstable state [80]. Intriguingly, the uuuduuud state is found to be stabilized by quantum fluctuations in the regime $h\gtrsim h_{\rm sat}/2$, whereas the stabilization by thermal fluctuations takes place at lower fields $h\lesssim h_{\rm sat}/2$. Such a trend is also known for the Heisenberg antiferromagnet on the triangular lattice, see, e.g., discussion in Ref. [102]. We expect that the supersolid phases found in the classical case below and above the uuuduuud phase exist also for $s=1/2$. While this expectation has been formulated earlier [74], a detailed investigation goes beyond the scope of the present work. Nevertheless, we mention that Ref. [81] found several non-trivial ordered states in the magnetization curves of the $s=1/2$ FSAFM, although it is not entirely clear to us if their Y-like and V-like phases match our 2:1:1 and 3:1 states below and above the plateau, respectively.
@@ Acknowledgements.
Johannes Richter was a prolific scientist. Regrettably, he did not witness completion of all of his projects. This is one of those that remained unfinished. Johannes Richter contributed the zero-temperature Lanczos calculations, part of the finite-temperature Lanczos results, and participated in the early stages of preparing the manuscript. We are grateful to Jörg Schulenburg for providing spinpack, the code with which many of the computations were performed, and to Jürgen Schnack for sharing his insights on the finite-temperature Lanczos method. The $N=36$ TPQ computations were performed with a precursor of XDiag [134] using sublattice coding [135].

### v2

@@ III.2.1 Specific heat
The first case $h/J_{1}=4$ lies just below the saturation field that is $h_{\rm sat}/J_{1}=4.2$ for $J_{2}/J_{1}=0.55$ (see Eq. (7)). The result for the $N=36$ system is shown in Fig. 11(a) and serves as a benchmark. The proximity to saturation permits us to perform exact diagonalization (ED) that covers the complete spectrum up to sufficiently high energies to yield essentially exact results for the entire temperature range shown. This ED curve is free from the statistical errors that are inherent to FTLM and thus serves as a reference. We see that the FTLM curve is very close to the ED one, and also the TPQ curve deviates from these two only by an amount that is consistent with the estimated error. The features observed in Fig. 11(a) around $T/J_{1}=0.01$ and $0.05$ are most likely due to finite-size effects reflecting the finite system size $N=36$ and thus have no physical meaning.
@@
According to Subec. III.1.2, there should be no $m=1/2$ plateau at $T=0$ for $J_{2}/J_{1}=0.5$. If the case $s=1/2$ is analogous to the classical case shown in Fig. 2, the maximum of $C$ at $T/J_{1}\approx 0.1$ for $1\lesssim h/J_{1}\lesssim 2$ observed in Fig. 12(a) might nevertheless indicate a transition out of an uuuduuud phase stabilized by thermal fluctuations. For $J_{2}/J_{1}=0.55$, we find an $m=1/2$ plateau in the range of fields $2.24\lesssim h/J_{1}\lesssim 2.60$, see Fig. 8(c). The maximum of $C$ for $T/J_{1}\lesssim 0.1$ that one observes in this range of fields in Fig. 12(b) probably corresponds to a transition out of this plateau state, as we have discussed in the context of Fig. 11(b). Invoking again the analogy with the classical case Fig. 2, the continuation of this maximum to lower fields $1.2\lesssim h/J_{1}\lesssim 2.2$ and at $T/J_{1}\lesssim 0.08$ might again correspond to a transition out of an uuuduuud state stabilized by thermal fluctuations. This suggests an intriguing scenario where the crossover from stabilization by quantum to thermal fluctuations around $h/J_{1}\approx 0.55$ would be accompanied by a minimum in the transition temperature.
In the low-field limit, the specific heat of Fig. 12(a,b) exhibits a maximum around $T/J_{1}\approx 0.2$. This could be just a fluctuation maximum. However, it could also be a signature of a finite-temperature phase transition for a plaquette or columnar valence-bond solid ground state as suggested, e.g., in Refs. [33, 41, 52, 55, 59]. Such a state would break lattice symmetries (translations and/or rotations) and could thus support a finite-temperature phase transition. However, the features observed at $h=0$ in Fig. 12(a,b) are broad. If they correspond to phase transitions, finite-size effects would thus be large. In any case, clarification of the exact meaning of these low-field low-temperature maxima in the specific heat is a question that goes beyond the scope of the present work.
@@ III.2.2 Entropy and magnetocaloric effect
Geometric frustration is expected to enhance the magnetocaloric effect [135, 136, 137, 134]. In the present case, we expect such an enhancement specifically at $J_{2}/J_{1}=1/2$ where the single-magnon spectrum becomes degenerate, recall Eq. (9). Such an enhancement was indeed observed in previous works [86, 87, 91] that scanned the ratio $J_{2}/J_{1}$ at a fixed temperature. Here, such an enhancement is observed as a pronounced minimum in the lowest three isentropes $T(h)$ around $h_{\rm sat}$ in Fig. 12(c).
@@ IV Conclusion
We have revisited the behavior of the $J_{1}$-$J_{2}$ FSAFM in a magnetic field, pushing previous investigations [72, 73, 74] to larger system sizes and higher accuracy.
@@
Then we turned to the spin-1/2 case. First, we presented zero-temperature magnetization curves for several ratios of $J_{2}/J_{1}$ and provided evidence for the emergence of a distinct $m=1/2$ plateau that can be understood as an uuuduuud state stabilized by quantum fluctuations. We find a relatively large stabilization regime $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$ that is shifted to values $J_{2}>J_{1}/2$ such that analytic treatment of this uuuduuud $m=1/2$ plateau requires expansion around a classically unstable state [80]. Intriguingly, the uuuduuud state is found to be stabilized by quantum fluctuations in the regime $h\gtrsim h_{\rm sat}/2$, whereas the stabilization by thermal fluctuations takes place at lower fields $h\lesssim h_{\rm sat}/2$. Such a trend is also known for the Heisenberg antiferromagnet on the triangular lattice, see, e.g., discussion in Ref. [104]. We expect that the supersolid phases found in the classical case below and above the uuuduuud phase exist also for $s=1/2$. While this expectation has been formulated earlier [74], a detailed investigation goes beyond the scope of the present work. Nevertheless, we mention that Ref. [81] found several non-trivial ordered states in the magnetization curves of the $s=1/2$ FSAFM, although it is not entirely clear to us if their Y-like and V-like phases match our 2:1:1 and 3:1 states below and above the plateau, respectively. Another point that might merit further attention is the question of the existence of the $m=1/3$ plateau predicted by a Chern-Simons theory [76]. A systematic numerical investigation would require a sequence of system sizes divisible by three, but here we have only two of those at our disposal, namely $N=36$ and $48$.
@@ Acknowledgements.
Johannes Richter was a prolific scientist. Regrettably, he did not witness completion of all of his projects. This is one of those that remained unfinished. Johannes Richter contributed the zero-temperature Lanczos calculations, part of the finite-temperature Lanczos results, and participated in the early stages of preparing the manuscript. We are grateful to Jörg Schulenburg for providing spinpack [138], the code with which many of the computations were performed, and to Jürgen Schnack for sharing his insights on the finite-temperature Lanczos method. The $N=36$ TPQ computations were performed with a precursor of XDiag [139] using sublattice coding [140].